OpenStudy (anonymous):

Find the traces for the multivariable equation y=z^2

OpenStudy (anonymous):

I am mainly confused about how you know which variable to make the trace equal to

OpenStudy (loser66):

The trace of a surface in a plane is the intersection of the surface with that plane. While we can discuss traces in any plane, for surfaces in the form x = f(y,z) we are particularly interested in traces in planes parallel to the yz plane. Note that the yz plane has the equation x = 0, and planes parallel to the yz plane have equations in the form x = a. To obtain the trace of x=f(y,z) in the plane x= a, we set f(y,z) =a hence x = f(y, z) = y -z^2 the trace in the plane x =1 is y - z^2 =1 the trace in the plane x =2 is y -z^2 =2 so far and so on.

OpenStudy (anonymous):

Whoa. f(y,z)? I'm not actually in multivariable. We are just doing an intro to that stuff in case any of us decides to to that, and we have not done any of that notation

OpenStudy (anonymous):

And we are only dealing with xy,xz, and yz btw

OpenStudy (loser66):

If you have z = f(x,y) , why can't you have x = f(y,z)??

OpenStudy (loser66):

Obviously, in 3 D , the equation y = z^2 implies x =0, right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So the xy trace is y=0?

OpenStudy (loser66):

hihihi... that's what I thought. No need to follow if you don't study it yet.

OpenStudy (anonymous):

ok

OpenStudy (loser66):

xy trace is z =0

OpenStudy (anonymous):

Why can't it be y=0

OpenStudy (loser66):

if the function is x,z, then y =0

OpenStudy (anonymous):

The function is y=z^2

OpenStudy (anonymous):

And it looks like this

OpenStudy (loser66):

that is y, z , since y = z^2 \(\implies\) y -z^2 =0

OpenStudy (anonymous):